An object has 255 kJ of potential energy and a mass of 4517 mg. Determine the object's height above Earth's surface.
1
Expert's answer
2018-03-13T08:53:32-0400
Solution:
E = mgh
h = E/(mg) = 255000/(0.004517·9.8) = 5.76·10 ^6 (m) = 5760 km
This is an unrealistic value, if there is a typo in the question and “mg” is “kg” the answer would
be 5.76 meters.
Answer: 5760 km